OpenStudy (anonymous):

How many different pizza's can be ordered if the restaurant offers 15 different toppings and there is not limit to the number of toppings on the pizza? Okay, so I know that the answer is 32768. But I'm just not positive how to get to that, could someone please explain?

OpenStudy (anonymous):

pizzas *

OpenStudy (hba):

k wait a sec

OpenStudy (anonymous):

yup

OpenStudy (hba):

sorry couldnt get it

OpenStudy (hba):

@Compassionate help needed

OpenStudy (anonymous):

Ther is no upper limit means it could go upto infinity..

OpenStudy (anonymous):

Otherwise I guess it's a combination problem.

OpenStudy (compassionate):

15 different toppings. 15(bits) \[(2^5)^3 = 32^3 = 32,768\]

OpenStudy (anonymous):

u can also do it this way: [although its reeeally lengthy!] 15C1 + 15C2 + 15C3 +........+ 15C14 + 16C15

OpenStudy (anonymous):

sorry, the last term should be 15C15

OpenStudy (anonymous):

thanks :D !!!